OpenStudy (anonymous):

A 15-ft diagonal brace on a bridge connects a support of the center of the bridge to a side support on the bridge. The horizontal distance that it spans is 3 ft longer than the height that it reaches on the side of the bridge. Find the horizontal and vertical distances spanned by this brace.

OpenStudy (mayankdevnani):

the horizontal distance ----> x the height -----> y The horizontal distance is 3 ft longer than the height -----> x = y + 3 Lets apply the Pythagorean Theorem: 15^2 = x^2 + y^2 225 = x^2 + (x - 3)^2 225 = x^2 + (x^2 - 6x + 9) 225 = 2x^2 - 6x + 9 0 = 2x^2 - 6x + 9 - 225 0 = 2x^2 - 6x - 216 ----> divide all by 2 0 = x^2 - 3x - 108 -----> FOIL 0 = (x - 12)(x + 9) x = 12 ( true ) . . .. x = -9 ( impossible negative numbers in length ) x = y + 3 ----> when x = 12. . .. . .. it implies y = 9

OpenStudy (mayankdevnani):

ok @sharron482

OpenStudy (anonymous):

ok

OpenStudy (mayankdevnani):

yaa

OpenStudy (mayankdevnani):

am i right? @ajprincess

OpenStudy (mayankdevnani):

am i right? @ganeshie8

OpenStudy (anonymous):

Yeah